OpenStudy (anonymous):

solve the equation. check your solution. 4=sqrt(h)-6

OpenStudy (anonymous):

\[4=\sqrt{h}-6\]

OpenStudy (compassionate):

Do you know the first step, hun?

OpenStudy (anonymous):

no, I looked at the book but I am completely lost, I cant follow their steps. I was thinking you need to add 6 to both sides first making it \[10=\sqrt{h}\] am I right?

OpenStudy (compassionate):

Yes. Go on.

OpenStudy (anonymous):

then you solve for sqrt to get h?

OpenStudy (compassionate):

There is no sqrt of H, since it's not h^2. So you perform the opposite action of a square root.

OpenStudy (anonymous):

don't understand the opposite action, think this is where I get lost

OpenStudy (compassionate):

h = 100, by the way.

OpenStudy (compassionate):

Want me to explain it to you?

OpenStudy (anonymous):

please in English not math, lol

OpenStudy (compassionate):

Alright. One second. Let me type something up to help you understand.

OpenStudy (anonymous):

k

OpenStudy (compassionate):

So we have the square root in the middle. How we undo that is by squaring both sides. So if we square both sides to get the x out, we also square the 10, which gives us 100. Make sense?

OpenStudy (anonymous):

so squaring 10 is 3.16, and h is 100

OpenStudy (anonymous):

wait that is not right is it?

OpenStudy (anonymous):

ow wait, square of h = 100, and if we take 10*10 that also =100

OpenStudy (anonymous):

lightbulb just went off.....squaring is opposite of square root!!!!!

OpenStudy (compassionate):

\[10 = \sqrt{h}\rightarrow 10^2 = \sqrt{h^2}\] Okay, so we know that if we have 2 terms under the square root symbol we can remov ethem. \[10^2 = \sqrt{h^2} = 100 = h\]

OpenStudy (compassionate):

CORRECT!

OpenStudy (anonymous):

\[4=\sqrt{100}-6\]

OpenStudy (compassionate):

So the same thing. 10 = sqrt{100} Now square both numbers 100 = 10,000 Which is false.

OpenStudy (anonymous):

ok, I see what you are saying....got it...I am going to write these steps down for the next problems, thanks and I appreciate your patience!!!!

OpenStudy (compassionate):

No problem. Remember to give me a medal !!

OpenStudy (anonymous):

;)